PHENOMENA TAKING PLACE AT THE ATOMIC LEVEL UNDER CERTAIN PERTURBATIONS

\underline{KHOLMURAD KHASANOV}, Samarkand University, 15, University Boulevard, Samarkand, Uzbekistan, 703004 (e-mail:hanvar@samarkand.uz).

In the talk^{a} the experimental results on a change of the gravitation mass of the substance under an influence of the electromagnetic field was discussed. In the work^{b} results on a change of the substance weight and antigravitation quantums directed from Earth’s surface upwards was given. In this report some theoretical interpretation of experimental results, mentioned above is presented: • If the atom or atomic system under external perturbation absorpt the energy then ”within atomic antigravitation” is observed, here takes place the decreasing of the atomic weight of the substance. • If the atom or atomic system radiate the energy then takes place ”within atomic gravitation”, then atomic weight of substance is increased. Observed phenomena ”within atomic gravitation and antigravitation”, in our opinion, is of large theoretical interest for wide discussion.

aKholmurad Khasanov. The Phenomena of the Electro-Gravitation. Abstracts of the Symposium on Experemental Gravitation. Samarkand, August 16-21, 1999.

bKholmurad Khasanov. Antigravitation Quantum High Energy, Abstracts of 55^{th} Ohio State University International Symposium on Molecular spectroscopy, June 12-16, 2000, T107, p.149.